# Virtual states and exponential decay in small-scale dynamo

### v2

@@ Abstract
We also find the critical Reynolds number and the increment of growth/decay above and under the threshold; we express them in terms of the quantitative characteristic properties of the velocity correlator, which makes it possible to compare the results with the data of different simulations.
Small-scale dynamo (SSD) in a turbulent flow is a widely discussed possible mechanism of magnetic field generation in various astrophysical objects Parker ; BSS-rev ; HazKit . The mechanism is based on stretching of magnetic lines trapped and carried by a small-scale turbulent flow, and implies high Reynolds numbers.
@@
either much larger or much smaller than 1. The first case corresponds to, e.g., interstellar and intergalactic matter, while the second takes place in the Sun ($Pm\sim 10^{-6}\div 10^{-4}$; Stix ; Rempel ) and solar-type stars, and inside planets. The question whether the SSD plays an important role in solar magnetic field generation remains open, in particular because of significant difficulties in modeling low Prandtl numbers in experiments and simulations Sch07 ; Hotta . However, during last decade, direct numerical simulations were performed for magnetic Prandtl numbers down to $10^{-1}$ and even lower Sch07 ; I ; B ; W .
The theoretical investigations of small-scale dynamo are based on the Kazantsev approach Kaz . The Kazantsev equation relates the second order one-time magnetic field correlator and the longitudinal velocity correlator of some special type. For low Prandtl numbers, this equation was analyzed in, e.g., NRS ; VK ; Verg ; Rog97 ; Vinc ; Bold ; finny ; Malyshkin ; Kleeorin12 ; Schob ; Yu ; mnras ; the critical magnetic Reynolds number
@@
corresponding to the threshold of generation was found NRS , and for $Rm>Rm_{c}$, the magnetic field growth rate near the threshold was shown to be Kleeorin12
@@
However, for $Rm$ below the generation threshold, from the Kazantsev theory it follows that the magnetic field perturbations spectrum is continuous; thus, the decay is not exponential, and $\left.\gamma\right|_{Rm<Rm_{c}}=0$. On the contrary, numerical simulations Sch07 ; I ; B ; W show that the linear dependence (1) still holds for $Rm<Rm_{c}$: in the graph $\gamma(\ln Rm)$ they present, the straight line continues under the abscissa axis.
@@ I Kazantsev equation
In this paper, following Kaz and many other papers, we consider Gaussian and delta time-correlated velocity statistics. Then, it is completely determined by the Lagrangian relative diffusivity of particles second order correlator:
@@
We are interested in the pair correlation function $\left\langle B_{i}(\mathbf{r},t)B_{j}(\mathbf{r}^{\prime},t)\right\rangle$. To get the corresponding equation, one has to express its time derivative by means of (2), and then average over different realizations of the velocity field. The cross correlations of magnetic induction and velocity can be split by means of the Furutsu-Novikov theorem,
@@
The conditions of isotropy and non-divergence of the velocity field reduce the tensor equation to only one component:
@@
and by means of the Laplace transform
$\psi=\int\limits_{0}^{\infty}e^{\gamma t}\psi_{\gamma}(\rho)d\gamma$ (8)
@@
Here and below we omit the index $\gamma$ in $\psi_{\gamma}(x)$. In what follows we specify the explicit shape of the potential, and consider the existence of solutions to this equation. Formally, this equation is Schrodinger equation with variable mass; to proceed to the Sturm-Liouville problem, one has to make one more change of variables and to remove the dependence on $\rho$ in the term containing $\gamma$. For $Pr_{m}\gg 1$ this is done in (apj, , Appendix C). However, qualitative behavior of the solution in the case of variable mass remains the same, so here we analyze it without additional transformations.
@@ II Definition of parameters
In a turbulent flow, the structure function $b(\rho)$ demonstrates roughly piecewise power behavior; one can describe it by extending the Vainshtein-Kichatinov model VK :
@@
Here the first line corresponds to the viscous range restricted by the Kolmogorov scale $r_{\nu}$, the second line describes the inertial range, and the third line demonstrates schematically the transition to the large-eddy scale where the power law breaks. Following Kazantsev, we leave the power $s$ undefined for now; the Kolmogorov theory predicts $s=1/3$ VK ; Rog97 , from intermittency it follows that $s>1/3$ Vain ; Frisch ; from the theoretical bridge relations LvovX combined with recent DNS results Iyer it follows that $s=0.39$ for asymptotically large Reynolds numbers.
The graph of this function in logarithmic coordinates has three linear sections; in real turbulence, of course, they are separated by two transition zones. The first transition zone between viscous and inertial ranges is called bottleneck; it does not affect generation seriously since for very small magnetic Prandtl numbers the magnetic diffusivity scale lies deep inside the inertial range. So, we are interested in the scales $\rho\gg r_{\nu}$, and thus can neglect the effects of viscosity and the bottleneck (i.e., the transition from viscous to inertial range). Instead, we concentrate on the inertial range and the transition to the integral scale. The second transition zone is essential for magnetic field generation paper140 , but to understand the main idea of the effect we do not have to consider it accurately.
@@
$Rm\simeq\frac{3b_{\infty}}{2\eta}=3\cdot X^{1+s}$
However, this is only a rough estimate to the order of magnitude, since different simulations use different definitions of $L$. For more accurate comparison of $X$ and $Rm$ , one has to investigate the properties of $b(r)$ used in every particular simulation. An attempt of such an accurate analysis will be made in our paper paper140 .
@@ IV Critical Reynolds number
The phase $\phi$ is determined by the behavior of the solution at small $x$. To find it, one has to consider the exact solution of (11) for $x<X$ that converges at small $x$ :
$\,\,\,\,\,\,\,\,\,\alpha(s)=\frac{4+s}{2(1+s)}\,\,\,\,\,\,\beta(s)=\frac{\kappa}{(1+s)}$
Comparing this with (15) and making use of
one can find
@@
This is the value of $X$ that corresponds to the generation threshold; other solutions correspond to appearance of next levels in the potential well. The corresponding critical value of $Rm$ is
@@
This estimate depends significantly on the exact definition of the Reynolds number. For detailed comparison of this theoretical result with the results of numerical simulations, see paper140 .
Without the last term in the potential, the matching conditions would take the form $\psi_{<}=\psi_{>}$, $\psi^{\prime}_{<}=\psi^{\prime}_{>}$, and one would get $\kappa\tan\left(\kappa\ln X+\varphi\right)=3/2$ instead of (16). This corresponds to the potential obtained in Kleeorin12 ; Schob and results in $X_{c}({\rm without\,\delta})=17.6$, so the value of $Rm_{c}$ calculated without the account of $\delta$-function in the last term is underestimated to about 20% .
@@ V.1 Generation: $\gamma>0$
So, it is convenient to present the matching conditions (13) in the form
Substituting (21) in the second equality, we get
Finally, the substitution of the explicit form of $\kappa$ results in
@@
By chance, for $s=1/3$, this coefficient is to good accuracy equal to 1. So, we obtain the coefficient in the logarithmic dependence of $\gamma$ on $Rm$. By means of the method proposed in (blobs, , Appendix A), we confirm this result from Eq. (11) numerically. To obtain the coefficient from DNS one needs the value of $r_{d}$. For accurate analysis of data of different papers and for their comparison with this result, we refer the reader to paper140 . Here we restrict ourselves by very rough estimate to the order of magnitude, to make sure that the theory does not contradict experimental/simulation data.
In B ; W the corresponding relation is presented in the form
@@ V.2 Damping: $\gamma<0$
Conside some fixed $X<X_{c}$, $X_{c}-X\ll X_{c}$. There are no solutions with positive $\gamma$, so the magnetic field correlations decay. The solution to Eq. (11) to the right from $X$ ($x>X$) is now described by a two-parametric oscillating function, so one can match the left-hand and right-hand sides for any $\gamma$. In terms of quantum mechanics, this continuous spectrum corresponds to positive energy of free particles. Since solutions with all $\gamma<0$ are present, the decay is slower than any exponential. (Also in two-dimensional turbulence, the decay is always power-law Kol ).
Thus, the Kazantsev theory predicts power-law decay for magnetic Reynolds numbers under the generation threshold. However, it turns out that just below the threshold, for each $X$ there is an exponent $\gamma(X)$ such that its amplitude in the inner region is much higher than the amplitudes corresponding to other exponents. This effect is analogous to a virtual level in quantum mechanics, and results in existence of temporal exponential damping with the exponent $\gamma(X)$. After some time, the exponential behavior changes back to a power-law.
@@
What is known about the ’outer’ amplitude $B$? It is, generally, a function of $q$ determined by the expansion of the initial function $\psi(t,x)$ in a series of the eigenfunctions (26) with different $q$. Assume that (unlike the case considered in Yu ) the initial function $\psi(t=0,x)$ is a function with a compact support $\ll X$. Then, according to (8), $\int\psi(x,q)dq=0$ for any $x$ larger than some $x_{1}$; the same is true for $f=\psi_{>}+x\psi_{>}^{\prime}$. From (27) we then get
@@
This estimate of the minimal $Rm$ that implies a virtual level is in concordance with the one that can be obtained from the height of the potential right after the peak (the delta function itself does not contribute to the appearance of virtual levels, only the ’thick’ part of the peak matters). Assuming that the boundary ’energy level’ equals to this height, we get from (11)
@@ VI Conclusion
Based on the Kazantsev theory, we show that the linear dependence $\gamma\propto\ln Rm/Rm_{c}$ continues also for negative argument, in full agreement with the results of the DNS Sch07 ; I ; B ; W . We predict that in longer observation, the exponential decay will stop after some time and be replaced by a slower decrease. We give an estimate of this time (34) and show that it is asymptotically long for magnetic Reynolds numbers close to the generation threshold, and by the end of the exponential stage the magnetic field correlator would decrease by $\exp{|\gamma|/\Delta\gamma}\sim\exp{|\gamma|^{-1/2}}$ times (33). For Reynolds magnetic numbers far from the critical value, approximately $Rm\lesssim Rm_{c}/2$, there is no stage of exponential decay.
Although the results were obtained in the frame of the extended Vainstein-Kichatinov model, the existence of the virtual level and, accordingly, the temporal exponential decay of the magnetic field correlator just below the generation threshold is a consequence of the flattening of the velocity correlation function at large scales, and hence, is a general feature of all reasonable models. It also seems that finite correlation time of the velocity field, its time irreversibility or helical asymmetry would not affect the result.

### v3

@@ Abstract
We also find the critical Reynolds number, the increment of growth/decay above and under the threshold and the time of exponential decrease under the threshold; we express them in terms of the quantitative characteristic properties of the velocity correlator, which makes it possible to compare the results with the data of different simulations.
Small-scale dynamo (SSD) in a turbulent flow is a widely discussed possible mechanism of magnetic field generation in various astrophysical objects Parker (1979); Brandenburg et al. (2012); Hazra et al. (2023). The mechanism is based on stretching of magnetic lines trapped and carried by a small-scale turbulent flow, and implies high Reynolds numbers.
@@
either much larger or much smaller than 1. The first case corresponds to, e.g., interstellar and intergalactic matter, while the second takes place in the Sun ($Pm\sim 10^{-6}\div 10^{-4}$Stix (2012); Rempel et al. (2023)) and solar-type stars, and inside planets. The question whether the SSD plays an important role in solar magnetic field generation remains open, in particular because of significant difficulties in modeling low Prandtl numbers in experiments and simulations Schekochihin et al. (2007); Hotta et al. (2016). However, during last decade, direct numerical simulations were performed for magnetic Prandtl numbers down to $10^{-1}$ and even lower Schekochihin et al. (2007); Iskakov et al. (2007); Brandenburg et al. (2018); Warnecke et al. (2023).
The theoretical investigations of small-scale dynamo are based on the Kazantsev approach Kazantsev (1968). The Kazantsev equation relates the second order one-time magnetic field correlator and the longitudinal velocity correlator of some special type. For low Prandtl numbers, this equation was analyzed in, e.g., Zeldovich et al. (1990); Vainshtein and Kichatinov (1986); Vergassola (1996); Rogachevskii and Kleeorin (1997); Vincenzi (2002); Boldyrev and Cattaneo (2004); Arponen and Horvai (2007); Malyshkin and Boldyrev (2010); Kleeorin and Rogachevskii (2012); Schober et al. (2012); Makarova et al. (2025); Kopyev et al. (2024); the critical magnetic Reynolds number
@@
corresponding to the threshold of generation was found Novikov et al. (1983), and for $Rm>Rm_{c}$, the magnetic field growth rate near the threshold was predicted in Rogachevskii and Kleeorin (1997) and rigorously derived in Kleeorin and Rogachevskii (2012)
@@
However, for $Rm$ below the generation threshold, from the Kazantsev theory it follows that the magnetic field perturbations spectrum is continuous; thus, the decay is not exponential, and $\left.\gamma\right|_{Rm<Rm_{c}}=0$. On the contrary, numerical simulations Schekochihin et al. (2007); Iskakov et al. (2007); Brandenburg et al. (2018); Warnecke et al. (2023) show that the linear dependence (1) still holds for $Rm<Rm_{c}$: in the graph $\gamma(\ln Rm)$ they present, the straight line continues under the abscissa axis.
@@ I Kazantsev equation
In this paper, following Kazantsev (1968) and many other papers, we consider Gaussian and delta time-correlated velocity statistics. Then, it is completely determined by the Lagrangian relative diffusivity of particles second order correlator:
@@
We are interested in the pair correlation function $\left\langle B_{i}(\mathbf{r},t)B_{j}(\mathbf{r}^{\prime},t)\right\rangle$. To get the corresponding equation, one has to express its time derivative by means of (2), and then average over different realizations of the velocity field. The cross correlations of magnetic induction and velocity can be split by means of the Furutsu-Novikov theorem Furutsu (1963); Novikov (1965),
@@
where $g[\mathbf{v}]$ is an arbitrary functional of $\mathbf{v}$. The conditions of isotropy and non-divergence of the velocity field reduce the tensor equation to only one component:
@@
and by means of the Laplace transform (after subtracting discrete positive modes)
$\psi=\int\limits_{-\infty}^{0}e^{\gamma t}\psi_{\gamma}(\rho)d\gamma$ (8)
@@
Here and below we omit the index $\gamma$ in $\psi_{\gamma}(x)$. In what follows we specify the explicit shape of the potential, and consider the existence of solutions to this equation. Formally, this equation is Schrodinger equation with variable mass; to proceed to the Sturm-Liouville problem, one has to make one more change of variables and to remove the dependence on $\rho$ in the term containing $\gamma$. For $Pr_{m}\gg 1$ this is done in (Kopyev et al., 2022, Appendix C). However, qualitative behavior of the solution in the case of variable mass remains the same, so here we analyze it without additional transformations.
@@ II Definition of parameters
In a turbulent flow, the structure function $b(\rho)$ demonstrates roughly piecewise power behavior; one can describe it by extending the Vainshtein-Kichatinov model Vainshtein and Kichatinov (1986):
@@
Here the first line corresponds to the viscous range restricted by the Kolmogorov scale $r_{\nu}$, the second line describes the inertial range, and the third line demonstrates schematically the transition to the large-eddy scale where the power law breaks. Following Kazantsev, we leave the power $s$ undefined for now; the Kolmogorov theory predicts $s=1/3$ Vainshtein and Kichatinov (1986); Rogachevskii and Kleeorin (1997), from intermittency it follows that $s>1/3$ Vainshtein (1982); Frisch (1995); from the theoretical bridge relations L’vov et al. (1997) combined with recent DNS results Iyer et al. (2020) it follows that $s=0.39$ for asymptotically large Reynolds numbers.
The graph of this function in logarithmic coordinates has three linear sections; in real turbulence, of course, they are separated by two transition zones. The first transition zone between viscous and inertial ranges is called bottleneck; it does not affect generation seriously since for very small magnetic Prandtl numbers the magnetic diffusivity scale lies deep inside the inertial range. So, we are interested in the scales $\rho\gg r_{\nu}$, and thus can neglect the effects of viscosity and the bottleneck (i.e., the transition from viscous to inertial range). Instead, we concentrate on the inertial range and the transition to the integral scale. The second transition zone is essential for magnetic field generation Kopyev et al. (2026), but to understand the main idea of the effect we do not have to consider it accurately.
@@
$Rm\simeq\frac{3b_{\infty}}{2\eta}=3\cdot X^{1+s}$ (10)
However, this is only a rough estimate to the order of magnitude: actually, different simulations use different definitions of $L$, as it is associated to the pumping scale. Turbulence properties are not universal at these scales, and the resulting values of Rm may differ from (10). For more accurate comparison of $X$ and $Rm$ , one has to investigate the properties of $b(r)$ used in every particular simulation. An attempt of such an accurate analysis will be made in our paper Kopyev et al. (2026).
@@ IV Critical Reynolds number
The phase $\phi$ is determined by the behavior of the solution at small $x$. One can find (see Appendix A)
@@
This is the value of $X$ that corresponds to the generation threshold; other solutions correspond to appearance of next levels in the potential well. The corresponding critical value of $Rm$ estimated in (10) is
@@
Without the last term in the potential, the matching conditions would take the form $\psi_{<}=\psi_{>}$, $\psi^{\prime}_{<}=\psi^{\prime}_{>}$, and one would get $\kappa\tan\left(\kappa\ln X+\varphi\right)=3/2$ instead of (17). This corresponds to the potential obtained in Kleeorin and Rogachevskii (2012); Schober et al. (2012) and results in $X_{c}({\rm without\,\delta})=17.6$, so the value of $Rm_{c}$ calculated without the account of $\delta$-function in the last term is underestimated to about 20% .
@@ V.1 Generation: $\gamma>0$
From the mathcing conditions (14) one can get (see Appendix B)
@@
By chance, for $s=1/3$, this coefficient is to good accuracy equal to 1. So, we obtain the coefficient in the logarithmic dependence of $\gamma$ on $Rm$. By means of the method proposed in (Il’yn et al., 2021, Appendix A), we confirm this result from Eq. (12) numerically. To obtain the coefficient from DNS one needs the value of $r_{d}$. For accurate analysis of data of different papers and for their comparison with this result, we refer the reader to Kopyev et al. (2026). Here we restrict ourselves by very rough estimate to the order of magnitude, to make sure that the theory does not contradict experimental/simulation data.
In Brandenburg et al. (2018); Warnecke et al. (2023) the corresponding relation is presented in the form
@@ V.2 Damping: $\gamma<0$
Consider some fixed $X<X_{c}$, $X_{c}-X\ll X_{c}$. There are no solutions with positive $\gamma$, so the magnetic field correlations decay. For $\gamma<0$, the solution to Eq. (12) to the right from $X$ ($x>X$) is described by a two-parametric oscillating function of $x$, so one can match the left-hand and right-hand sides for any $\gamma$. In terms of quantum mechanics, this continuous spectrum corresponds to positive energy of free particles. Since solutions with all $\gamma<0$ are present, the decay is slower than any exponential. (Also in two-dimensional turbulence, the decay is always power-law Novikov et al. (1983); Kolokolov (2017)).
Thus, the Kazantsev theory predicts power-law decay for magnetic Reynolds numbers under the generation threshold. However, it turns out that just below the threshold, for each $X$ there is an exponent $\gamma(X)$ such that the maximal value of $\psi_{\gamma}$ in the inner region is much higher than the corresponding values for other $\gamma$-s. In quantum mechanics this effect is usually interpreted as existence of virtual levels that make up a quasi-discrete spectrum Flügge (2012); Kleeorin and Rogachevskii (1994). For Eq. (12) the quasi-discrete spectrum results in existence of exponential damping with the exponent $\gamma(X)$. After some time, which we find in this section, the exponential behavior changes to a power-law.
@@
What is known about the ’outer’ amplitude $B$? It is, generally, a function of $q$ determined by the expansion of the initial function $\psi(t,x)$ in a series of the eigenfunctions (25) with different $q$. Assume that (unlike the case considered in Yushkov et al. (2018)) the initial function $\psi(t=0,x)$ is a function with a compact support $\ll X$. Then, according to (8), $\int\psi(x,q)dq=0$ for any $x$ larger than some $x_{1}$; the same is true for $f=\psi_{>}+x\psi_{>}^{\prime}$. From (26) we then get
@@
This estimate of the minimal $Rm$ that implies a virtual level is in concordance with the one that can be obtained from the height of the potential right after the peak Kleeorin and Rogachevskii (1994) (the delta function itself does not contribute to the appearance of virtual levels, only the ’thick’ part of the peak matters). Assuming that the boundary ’energy level’ equals to this height, we get from (12)
@@ VI Conclusion
Based on the Kazantsev theory, we show that the dependence $\gamma\propto\ln Rm/Rm_{c}$ continues also for $Rm<Rm_{c}$, in full agreement with the results of the DNS Schekochihin et al. (2007); Iskakov et al. (2007); Brandenburg et al. (2018); Warnecke et al. (2023). We predict that in longer observation, the exponential decay will stop after some time and be replaced by a slower decrease. We give an estimate of this time (33) and show that it is asymptotically long for magnetic Reynolds numbers close to the generation threshold, and by the end of the exponential stage the magnetic field correlator would decrease by $\exp{|\gamma|/\Delta\gamma}\sim\exp{|\gamma|^{-1/2}}$ times (32). For Reynolds magnetic numbers far from the critical value, approximately $Rm\lesssim Rm_{c}/2$, there is no stage of exponential decay.
Although the results were obtained in the frame of the extended Vainstein-Kichatinov model, the existence of the virtual level and, accordingly, the temporal exponential decay of the magnetic field correlator just below the generation threshold is a consequence of the flattening of the velocity correlation function at large scales, and hence, is a general feature of all reasonable models. Since the levels are determined by a Sturm-Liouville-type boundary problem, it also seems that finite correlation time of the velocity field, its time irreversibility or helical asymmetry would not affect the result.
@@
Appendix A Derivation of (18)
To find $\phi$, one has to consider the exact solution of (12) for $x<X$ that converges at small $x$ :
$\,\,\,\,\,\,\,\,\,\alpha(s)=\frac{4+s}{2(1+s)}\,\,\,\,\,\,\beta(s)=\frac{\kappa}{(1+s)}$
Comparing this with (16) and making use of
one can get (18)
Appendix B Derivation of (23)
It is convenient to present the matching conditions (14) in the form
Substituting (21) in the second equality, we get
Finally, the substitution of the explicit form of $\kappa$ results in (23).